OpenStudy (anonymous):

Solve: 7(x + 2) = 49

OpenStudy (anonymous):

x + 2 = 7 x = 5

OpenStudy (anonymous):

yep

OpenStudy (hoblos):

7x +14 = 49 x= 35/7 =5

OpenStudy (anonymous):

Thank's for explaining Hoblos.

OpenStudy (hoblos):

you are welcome ;)